INFRARED SPECTRA OF CO_{2} AND H_{2}O AT ATMOSPHERIC TEMPERATURES

\underline{CHENGBO SUN}, QUNJUN ZOU, YOU-KYOUNG CHUNG, PRASAD VARANASI, Institute for Terrestrial and Planetary Atmospheres State University of New York at Stony Brook Stony Brook, NY11794-5000 Email:pvaranasi@notes.cc.sunysb.edu.

We have measured the absolute intensities as well as N_{2}- and air- broadened half-widths of lines in the P- and R- branches of the v2- fundamental bands of CO_{2} at temperatures between 170K and 296K employing a Fourier-Transform-Spectrometer with spectral resolution as high as 0.0016 cm^{−1}. Intensities and self-broadening coefficients of lines of water vapor in the 7000-8800cm^{−1} region have also been measured at room temperature.